# Wiman-Valiron method for fractional derivatives and sharp growth estimates of $\alpha$-analytic solutions for linear fractional differential equations

### v1

@@ 4 $\alpha$-analyticity of soultions for (1)
$\frac{\Gamma((s+k)\alpha+1)}{\Gamma(s\alpha+1)}\leq{M_{2}}{((s+k)\alpha+1)^{k\alpha}}=:\delta_{s,k},\quad s,k\in\mathbb{Z}_{+},n\in\mathbb{N}.$ (8)
The values $\beta_{0}$, …, $\beta_{n-1}$ are chosen recursively so that the equality in (6) holds, i.e.
@@
where the constant $C$ will be specified later.
We prove (6) by induction. Since the induction base holds by the choice of $\beta_{0},\dots,\beta_{n-1}$, it is sufficient to prove the induction step.
Let $m\geq 0$, and (6) hold for $0\leq p\leq m+n-1$. Then by (4)
@@
$\leq\gamma_{m}M\sum_{k=0}^{n-1}\delta_{m,k}\sum_{s=0}^{m}\frac{\beta_{k+s}}{r^{\alpha(m-s)}}\prod_{j=1}^{k+s-1}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right).$ (9)
@@
$=\prod_{j=1}^{k-1}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right)\prod_{j=k}^{k+m}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right)=\prod_{j=1}^{k+m}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right)$
Substituting this estimate into (9) we get
@@
$\leq M_{1}M_{2}\sum_{k=0}^{n-1}\frac{((m+k)\alpha+1)^{k\alpha}}{(m\alpha+1)^{n\alpha}}\prod_{j=1}^{k+m}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right)$
@@
If all coefficients $P_{j}$ of (1) are polynomilals, then all $\alpha$-analytic soultions have the form $v(t^{\alpha})$ where $v$ is an entire function of finite order of the growth.
@@
By Theorem 2, there is an entire function $v$ such that $y(t)=v(t^{\alpha})$ is the unique solution to the Cauchy problem (1), (1). By Theorem A, there exists a set $E\subset[1,\infty)$ of finite logarithmic measure such that
$\mathcal{D}_{\alpha}^{j}v(z)=(\nu(r,v)\alpha)^{j\alpha}\frac{v(z)}{z^{j}}(1+o(1)),\quad|z|\not\in E,$ (10)
@@
Let $c_{j}z^{d_{j}}$ be the leading coefficient of $P_{j}(z)$, $j\in\{0,\dots,n-1\}$. Substituting (10) into (1) and dividing by $v(z)$ we obtain $(\nu=\nu(|z|,v))$
@@
Let $P_{j}$ be polynomials of degree $d_{j}=\deg P_{j}$, $j\in\{0,\dots,n-1\}$, $p_{0}\not\equiv 0$, and $\max_{0\leq k\leq n-1}\frac{d_{k}}{n-k}=\frac{d_{0}}{n}$. Then all non-trivial $\alpha$-analytic solutions $u$ of the equation (1) has the form $u(t)=f(t^{\alpha})$, $t\geq 0$, where the order of an entire function $f$ is $\rho(f)=\frac{1}{\alpha}\left(1+\frac{d_{0}}{n}\right)$.
@@
Suppose the contrary. Then, by (11) there exists $\eta>0$ and a sequence of positive numbers $(r_{n})$ tending to $\infty$ such that $1<r_{m}<r_{m+1}/2$ with $\nu(r_{m})\geq r_{m}^{\sigma_{0}+\eta}$. Let $F=\bigcup_{m=1}^{\infty}[r_{m},2r_{m}]$. Clearly, $F$ has infinite logarithmic measure. Moreover, for $r\in F$, we have that $r\in[r_{m},2r_{m}]$ for some $m=m(r)$. Since $\nu(r)$ is non-decreasing,
$\nu(r)\geq\nu(r_{m})\geq r_{m}^{\sigma_{0}+\eta}\geq\frac{r^{\sigma_{0}+\eta}}{2^{\sigma_{0}+\eta}},\quad r\in F.$ (12)
Therefore, for $r\in F\setminus E$, which is of infinite logarithmic measure, and, in particular, unbounded, we have that (11) holds. Note that for every $j\in\{1,\dots,n-1\}$ and $\varepsilon>0$ the following estimates are valid
@@
which contradicts (12) provided that $\varepsilon\in(0,\eta n)$. Thus, $\sigma(f)\leq\sigma_{0}$.
@@
where $\frac{d_{0}}{n}=\max_{0\leq k\leq n-1}\frac{d_{k}}{n-k}$ for any real $s<\sigma_{0}\alpha$. That is $\nu(r)=O(r^{\sigma})$, $\sigma<\sigma_{0}$ is also impossible, because in this case (11) can be rewritten in the form $(c_{0}+o(1))z^{d_{0}}=0$. The theorem is proved. ∎

### v2

@@ 3 Proof of Theorem 1
$\frac{\Gamma((s+k)\alpha+1)}{\Gamma(s\alpha+1)}\leq{M_{2}}{((s+k)\alpha+1)^{k\alpha}}=:\delta_{s,k},\quad s,k\in\mathbb{Z}_{+},n\in\mathbb{N}.$ (4.8)
The values $\beta_{0}$, …, $\beta_{n-1}$ are chosen recursively so that the equality in (4.6) holds, i.e.
@@
where the constant $C$ will be specified later. It is an elementary exercise to prove that $\beta_{p}\to 0$ as $p\to\infty$.
We prove (4.6) by induction. Since the induction base holds by the choice of $\beta_{0},\dots,\beta_{n-1}$, it is sufficient to prove the induction step.
Let $m\geq 0$, and (4.6) hold for $0\leq p\leq m+n-1$. Then by (4.4)
@@
$\leq\gamma_{m}M\sum_{k=0}^{n-1}\delta_{m,k}\sum_{s=0}^{m}\frac{\beta_{k+s}}{r^{\alpha(m-s)}}\prod_{j=1}^{k+s-1}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right).$ (4.9)
@@
$=\prod_{j=1}^{k-1}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right)\prod_{j=k}^{k+m}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right)=\prod_{j=1}^{k+m}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right).$
Substituting this estimate into (4.9) we get
@@
$\leq MM_{1}M_{2}\sum_{k=0}^{n-1}\frac{((m+k)\alpha+1)^{k\alpha}}{(m\alpha+1)^{n\alpha}}\prod_{j=1}^{k+m}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right)$
@@
If all coefficients $P_{j}$ of (1.1) are polynomilals, then all $\alpha$-analytic soultions have the form $v(t^{\alpha})$ where $v$ is an entire function of finite order of the growth.
@@
By Theorem 2, there is an entire function $v$ such that $y(t)=v(t^{\alpha})$ is the unique solution to the Cauchy problem (1.1), (4.1). By Theorem A, there exists a set $E\subset[1,\infty)$ of finite logarithmic measure such that
$\mathcal{D}_{\alpha}^{j}v(z)=(\nu(r,v)\alpha)^{j\alpha}\frac{v(z)}{z^{j}}(1+o(1)),\quad|z|\not\in E,$ (4.10)
@@
Let $c_{j}z^{d_{j}}$ be the leading coefficient of $P_{j}(z)$, $j\in\{0,\dots,n-1\}$. Substituting (4.10) into (1.1) and dividing by $v(z)$ we obtain $(\nu=\nu(|z|,v))$
@@
Let $P_{j}$ be polynomials of degree $d_{j}=\deg P_{j}$, $j\in\{0,\dots,n-1\}$, $p_{0}\not\equiv 0$, and $\max_{0\leq k\leq n-1}\frac{d_{k}}{n-k}=\frac{d_{0}}{n}$. Then all non-trivial $\alpha$-analytic solutions $y$ of the equation (1.1) has the form $y(t)=f(t^{\alpha})$, $t\geq 0$, where the order of an entire function $f$ is $\rho(f)=\frac{1}{\alpha}\left(1+\frac{d_{0}}{n}\right)$.
Corollary.
Let $\alpha>0$, $n\in\mathbb{N}$, $P$ be a nontrivial polynomial of degree $d_{0}$. All non-trivial $\alpha$-analytic solutions $y$ of the equation
$\mathbb{D}_{\alpha}^{n}y+P(x^{\alpha})y=0$
has the form $y(t)=f(t^{\alpha})$, $t\geq 0$, where the order of an entire function $f$ is $\rho(f)=\frac{1}{\alpha}\left(1+\frac{d_{0}}{n}\right)$.
@@
Suppose the contrary. Then, by (1.11) there exists $\eta>0$ and a sequence of positive numbers $(r_{n})$ tending to $\infty$ such that $1<r_{m}<r_{m+1}/2$ with $\nu(r_{m})\geq r_{m}^{\sigma_{0}+\eta}$. Let $F=\bigcup_{m=1}^{\infty}[r_{m},2r_{m}]$. Clearly, $F$ has infinite logarithmic measure. Moreover, for $r\in F$, we have that $r\in[r_{m},2r_{m}]$ for some $m=m(r)$. Since $\nu(r)$ is non-decreasing,
$\nu(r)\geq\nu(r_{m})\geq r_{m}^{\sigma_{0}+\eta}\geq\frac{r^{\sigma_{0}+\eta}}{2^{\sigma_{0}+\eta}},\quad r\in F.$ (4.12)
Therefore, for $r\in F\setminus E$, which is of infinite logarithmic measure, and, in particular, unbounded, we have that (4.11) holds. Note that for every $j\in\{1,\dots,n-1\}$ and $\varepsilon>0$ the following estimates are valid
@@
which contradicts (4.12) provided that $\varepsilon\in(0,\eta n)$. Thus, $\sigma(f)\leq\sigma_{0}$.
@@
where $\frac{d_{0}}{n}=\max_{0\leq k\leq n-1}\frac{d_{k}}{n-k}$ for any real $s<\sigma_{0}\alpha$. That is $\nu(r)=O(r^{\sigma})$, $\sigma<\sigma_{0}$ is also impossible, because in this case (4.11) can be rewritten in the form $(c_{0}+o(1))z^{d_{0}}=0$. The theorem is proved. ∎